# MORPH: Shape-agnostic PDE Foundation Models

### v2

@@ C.2 Autoregressive rollouts
Figs. 6 and 7 present 10-step autoregressive (AR) rollouts for MORPH-FM-Ti and MORPH-FM-S fine-tuned on the forced incompressible Navier–Stokes with Kolmogorov forcing (FNS-KF) dataset, initialized from the t=0 snapshot. Over the 10-step horizon, MORPH-FM-S achieves consistently lower error than MORPH-FM-Ti. Both models remain stable under rollouts with limited error accumulation and no concerning instability across the rollout.
Figure 6: Finetuning MORPH-FM-Ti ($\sim$ 7M) for FNS-KF prediction task: 10-step autoregressive rollouts with $t=0$ (initial frame) as input.
Figure 7: Finetuning MORPH-FM-S ($\sim$ 30M) for FNS-KF prediction task: 10-step autoregressive rollouts with $t=0$ (initial frame) as input.

### v3

@@ C.3.1 Standalone surrogate
C.3.2 Finetuned foundation models
Figs. 8 and 9 present 10-step autoregressive (AR) rollouts for MORPH-FM-Ti and MORPH-FM-S fine-tuned on the forced incompressible Navier–Stokes with Kolmogorov forcing (FNS-KF) dataset, initialized from the t=0 snapshot. Over the 10-step horizon, MORPH-FM-S achieves consistently lower error than MORPH-FM-Ti. Both models remain stable under rollouts with limited error accumulation and no concerning instability across the rollout.
Figure 8: Finetuning MORPH-FM-Ti ($\sim$ 7M) for FNS-KF prediction task: 10-step autoregressive rollouts for $v_{x}$ and $v_{y}$ with $t=0$ (initial frame) as input.
Figure 9: Finetuning MORPH-FM-S ($\sim$ 30M) for FNS-KF prediction task: 10-step autoregressive rollouts for $v_{x}$ with $t=0$ (initial frame) as input.
C.3.3 Finetuned foundation model (Pretrained with PDEGym)
NS-SL.
Fig. 10 presents full autoregressive (AR) rollouts for MORPH-FM-S*(27M), pretrained on PDEGym (NS-Sines, NS-Gaussians, CE-RP, CE-CRP, CE-KH, CE-Gauss) and subsequently fine-tuned on 128 trajectories of NS-SL and evaluated on 240 trajectories over 200 epochs. We demonstrate stable rollouts with limited error accumulation even with a compact 27M-parameter MORPH model. These results suggest that MORPH can serve as an efficient foundation model, accurately capturing the underlying dynamics while remaining computationally lightweight.
(a) First 10 time-steps of horizontal-velocity field ($v_{x}$).
(b) Last 10 time-steps of horizontal-velocity field ($v_{x}$).
(c) First 10 time-steps vertical-velocity field ($v_{y}$).
(d) Last 10 time-steps vertical-velocity field ($v_{y}$).
Figure 10: Full autoregressive rollouts for MORPH-FM-S*(27M) finetuned for NS-SL prediction task with 128 trajectories, tested on 240 trajectories for 200 epochs
NS-PwC.
Fig. 11 presents full autoregressive (AR) rollouts for MORPH-FM-S*(27M), pretrained on PDEGym (NS-Sines, NS-Gaussians, CE-RP, CE-CRP, CE-KH, CE-Gauss) and subsequently fine-tuned on 128 trajectories of NS-PwC and evaluated on 240 trajectories over 200 epochs. We obtain stable rollouts with limited error accumulation even using a compact 27M-parameter MORPH model. This indicates that MORPH can function as an efficient foundation model, capturing the essential dynamics while remaining computationally efficient.
(a) First 10 time-steps of horizontal-velocity field ($v_{x}$).
(b) Last 10 time-steps of horizontal-velocity field ($v_{x}$).
(c) First 10 time-steps vertical-velocity field ($v_{y}$).
(d) Last 10 time-steps vertical-velocity field ($v_{y}$).
Figure 11: Full autoregressive rollouts for MORPH-FM-S*(27M) finetuned for NS-PwC prediction task with 128 trajectories, tested on 240 trajectories for 200 epochs
CE-RM.
Fig. 12 presents full autoregressive (AR) rollouts for MORPH-FM-S*(27M), pretrained on PDEGym (NS-Sines, NS-Gaussians, CE-RP, CE-CRP, CE-KH, CE-Gauss) and subsequently fine-tuned on 128 trajectories of CE-RM and evaluated on 240 trajectories over 200 epochs. We observe stable rollouts, but note missing higher spectral components and finer details toward the end of the rollout. This behavior is expected given the relatively small model size (27M parameters) and the limited number of fine-tuning trajectories (128).
(a) First 10 time-steps of pressure field ($P$).
(b) Last 10 time-steps of of pressure field ($P$).
(c) First 10 time-steps of horizontal-velocity field ($v_{x}$).
(d) Last 10 time-steps of horizontal-velocity field ($v_{x}$) .
Figure 12: Full autoregressive rollouts for MORPH-FM-S*(27M) finetuned for CE-RM prediction task with 128 trajectories, tested on 240 trajectories for 200 epochs
Current limitations
MORPH is pretrained primarily on fluid-related datasets. However, a general-purpose PDE foundation model should ultimately encompass a broader range of physics, including elasticity problems such as fracture mechanics, wave propagation in solids, and fluid–structure interaction problems. At present, MORPH is restricted to a maximum of 500M parameters. With access to a larger and more diverse pretraining corpus, we expect that scaling the model to a few billion parameters will be necessary to fully exploit the available data. Moreover, the convolutional preprocessing and patching scheme is designed for regular grids, which limits direct applicability to unstructured meshes or irregular geometries. Finally, UPTF-7 employs scalar-to-vector lifting, which introduces additional batch-level memory and computational overhead.